### v2

# Hamiltonian reductions as affine closures of cotangent bundles

Abstract.
Let $Y$ be an irreducible non-singular affine $G$-variety with a $2$-large action. We show that the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities, isomorphic to the affine closure of $T^{*}Z_{\operatorname{reg}}$ where $Z\coloneqq Y/\!/G$. Furthermore, we provide sufficient conditions for the non-existence of a symplectic resolution for such varieties. These results yield three main applications: (i) providing a short proof of G. Schwarz’s theorem on the graded surjectivity of the push-forward map $\mathcal{D}(Y)^{G}\to\mathcal{D}(Z)$; (ii) establishing the surjectivity of the symbol map on $Z$; and (iii) confirming the non-linear analog of a conjecture of Kaledin–Lehn–Sorger for $2$-large actions.
Key words and phrases:
Hamiltonian reduction, symplectic singularity, Hamiltonian slice
2020 Mathematics Subject Classification:
14B05, 53D20, 14L24, 14J42
1. Introduction
Symplectic varieties, as introduced by Beauville [Bea00], are singular counterparts of hyperkähler manifolds and they are playing important roles not only in algebraic symplectic geometry, but also in geometric representation theory, 3d $\mathcal{N}=4$ supersymmetric gauge theory and so on.
Recently, we started a systematic investigation of a geometric construction of symplectic varieties through cotangent bundles. Namely for an irreducible non-singular variety $Y$ with $T^{*}Y$ its cotangent bundle, denote by $\mathcal{O}(T^{*}Y)$ the ring of regular functions on $T^{*}Y$. Then the affine closure $\mathcal{Z}_{Y}$ of $T^{*}Y$ is defined as the normal affine scheme ${\rm Spec}(\mathcal{O}(T^{*}Y))$ so that we have the natural affinization morphism
$\psi_{Y}\colon T^{*}Y\longrightarrow\mathcal{Z}_{Y}\coloneqq{\rm Spec}(\mathcal{O}(T^{*}Y)).$
When $Y$ is a non-singular projective variety, it was shown in [FL25b] that $\mathcal{Z}_{Y}$ is a symplectic variety if the tangent bundle $T_{Y}$ is big and if the projective bundle $\mathbb{P}T_{Y}$ is of Fano type. The most interesting case is where $Y$ is quasi-affine but not affine. In this case, $\psi_{Y}$ is just an open embedding and, therefore, $\mathcal{Z}_{Y}$ is the affine closure of $T^{*}Y$. It remains a challenging problem to determine when $\mathcal{Z}_{Y}$ is a symplectic variety.
Another geometric way to construct symplectic varieties is through Hamiltonian reductions, which is a powerful tool in both physics and geometry to produce new symplectic manifolds from a Hamiltonian action on a symplectic manifold. To be more precise, let $(X,\omega)$ be a non-singular affine symplectic variety on which there exists a Hamiltonian action by a reductive complex linear algebraic group $G$ so that the associated moment map $\mu:X\to\mathfrak{g}^{*}$ is $G$-equivariant (see § 2). We denote by $N_{X}$ the scheme-theoretic fiber $\mu^{-1}(0)$, called the shell. The Hamiltonian reduction X///GX/\!\!/\!\!/G is by definition the affine GIT quotient:
$N_{X}/\!/G=\mu^{-1}(0)/\!/G\coloneqq{\rm Spec}(\mathcal{O}({\mu^{-1}(0)})^{G}).$
Even for the case where $X$ is a symplectic linear representation of $G$, the Hamiltonian reduction X///GX/\!\!/\!\!/G can be non-reduced or reducible or non-normal.
If $Y$ is an irreducible non-singular affine $G$-variety, its cotangent bundle $T^{*}Y$ admits a natural Hamiltonian $G$-action. Let Y//GY/\!/G be the affine GIT quotient. Our first main result in this paper is to bridge the two geometric constructions as follows (see Definition 4.1 for $2$-large actions):
Theorem 1.1.
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Then the following holds:
  1. (1.1.1)
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities.
  2. (1.1.2)
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is the affine closure of $T^{*}(Y/\!/G)_{\rm reg}$. In particular, if Y//GY/\!/G is non-singular, then $T^{*}Y/\!\!/\!\!/G\cong T^{*}(Y/\!/G)$.
  3. (1.1.3)
    The Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ is non-singular if and only if Y//GY/\!/G is non-singular.
The $2$-largeness condition in Theorem 1.1 is not overly restrictive: as shown in [Sch95, Corollary 11.6 (2)] (see also [HSS20, Theorem 3.6]), most $G$-modules are in fact $2$-large, with only finitely many exceptions under mild assumptions. On the other hand, for the special case in which $Y$ is a linear representation of $G$, part (1.1.1) of Theorem 1.1 was proved in [HSS20] (see also [HSS24a, HSS24b] for other related results).
As an application of Theorem 1.1, we investigate the ring of differential operators on the quotient Y//GY/\!/G. To state our results precisely, we recall the following notation (see also § 5 for further background). For an irreducible normal variety $W$, let $\mathcal{D}(W)$ denote the ring of global differential operators on $W$, and define the graded $\mathbb{C}$-algebra
$S(W)\coloneq\bigoplus_{p\geq 0}H^{0}(W_{\operatorname{reg}},{\sf S}^{p}T_{W_{\operatorname{reg}}})=\mathcal{O}(T^{*}W_{\operatorname{reg}}).$
Theorem 1.2.
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Set $Z\coloneqq Y/\!/G$ with the quotient map $\pi_{Y}\colon Y\rightarrow Z$.
  1. (1.2.1)
    [Sch95, Theorem 0.4] The canonical morphism $(\pi_{Y})_{*}\colon\mathcal{D}(Y)^{G}\rightarrow\mathcal{D}(Z)$ is graded surjective.
  2. (1.2.2)
    The symbol map $\sigma_{Z}\colon\mathcal{D}(Z)\to S(Z)$ is surjective.
We next investigate the ($\mathbb{Q}$-)factoriality of the Hamiltonian reduction $T^{*}Y/\!\!/\!\!/G$ and use this to deduce the non‑existence of symplectic resolutions under certain conditions. Recall that the abelianization of $G$ is $G_{\operatorname{ab}}\coloneqq G/[G,G]$. A group $G$ is called perfect if $G_{\operatorname{ab}}=\{e\}$, i.e., $G=[G,G]$.
Theorem 1.3.
Let $G$ be a reductive group and $Y$ an irreducible non-singular affine $G$-variety that is $2$-large. Assume that $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$ or $G$ is connected.
  1. (1.3.1)
    If $G_{\operatorname{ab}}$ is finite and $\operatorname{Pic}(Y)$ is finite, then $\operatorname{Cl}(T^{*}Y/\!\!/\!\!/G)$ is finite. In particular, $T^{*}Y/\!\!/\!\!/G$ is $\mathbb{Q}$-factorial.
  2. (1.3.2)
    If $G$ is perfect and $\operatorname{Pic}(Y)$ is trivial, then $\operatorname{Cl}(T^{*}Y/\!\!/\!\!/G)$ is trivial; that is, $T^{*}Y/\!\!/\!\!/G$ is factorial.
We present a partial converse of Theorem 1.3 in Corollary 4.8. For $Y$ a $3$-large linear representation of $G$, the local factoriality and $\mathbb{Q}$-factoriality of $T^{*}Y/\!\!/\!\!/G$ were already established in [BS19, Theorem 1.2]. By combining Theorem 1.1 and Theorem 1.3, we obtain the following result.
Theorem 1.4.
Let $G$ be a semisimple group, and let $Y$ be an irreducible non-singular affine $G$-variety which is $2$-large. Assume that $\operatorname{Pic}(Y)$ is finite. Then $T^{*}Y/\!\!/\!\!/G$ admits a symplectic resolution if and only if Y//GY/\!/G is non-singular.
Theorem 1.4 yields a positive answer to a non-linear analog for all $2$-large actions of the following conjecture of D. Kaledin, M. Lehn and C. Sorger, which in the linear case follows from [BS19, Corollary 1.3].
Conjecture 1.5.
Let $G$ be a semisimple group and let $V$ be a finite dimensional representation of $G$. If there exists a symplectic resolution of $T^{*}V/\!\!/\!\!/G$, then V//GV/\!/G is non-singular.
Acknowledgments
Both authors are supported by the National Key Research and Development Program of China (No. 2025YFA1017302), the CAS Project for Young Scientists in Basic Research (No. YSBR-033) and the NSFC grant (No. 12288201). J. Liu is supported by the Youth Innovation Promotion Association CAS. We thank G. Bellamy and G. Schwarz for comments on a first version of this text.
2. Quasi-Hamiltonian varieties
Throughout this paper, we work over complex numbers, and $G$ always denotes a reductive complex linear algebraic group. By varieties we mean separated reduced schemes of finite type over $\mathbb{C}$, which are not necessarily irreducible.
2.1. Hamiltonian actions
Definition 2.1.
Let $X$ be an irreducible normal variety.
  1. (2.1.1)
    We call $X$ quasi-symplectic if there exists a regular symplectic form $\omega$ on its non-singular locus $X_{\operatorname{reg}}$.
  2. (2.1.2)
    We call $X$ symplectic if it is quasi-symplectic and for any resolution $f\colon Z\rightarrow X$, the form $f^{*}\omega$ extends to a regular $2$-form on $Z$.
Let $X$ be a non-singular symplectic variety equipped with an action of $G$ by symplectomorphisms. For a regular function $H\in\mathcal{O}(X)$, the skew-gradient $V_{H}$ of $H$ is, by definition, the regular vector field on $X$ given by the following equality:
$\omega(V_{H},\cdot)=dH.$
To any element $\xi\in\mathfrak{g}$ one associates the velocity vector field $V_{\xi}$.
Definition 2.2.
The action $G$ on $X$ equipped with a linear map $\mathfrak{g}\rightarrow\mathcal{O}(X)$, $\xi\rightarrow H_{\xi}$, is said to be Hamiltonian if it satisfies the following two conditions:
  1. (H1)
    The linear map $\mathfrak{g}\rightarrow\mathcal{O}(X)$ is $G$-equivariant.
  2. (H2)
    $V_{H_{\xi}}=V_{\xi}$.
Then $X$ is called a Hamiltonian $G$-variety so that $V_{\xi}$ and $H_{\xi}$ are called the Hamiltonian vector field and the Hamiltonian function associated with $\xi$, respectively.
For a Hamiltonian action $G$ on $X$, we can define a morphism $\mu_{G,X}\colon X\rightarrow\mathfrak{g}^{*}$ using the following formula:
$\langle\mu_{G,X}(x),\xi\rangle\coloneqq H_{\xi}(x),\quad\xi\in\mathfrak{g},\quad x\in X.$
This morphism is called the moment map of the Hamiltonian $G$-variety $X$ so that $\mu_{G,X}$ is $G$-equivariant with respect to the coadjoint action of $G$ on $\mathfrak{g}^{*}$ such that $H_{\xi}\in\mathcal{O}(X)$ is nothing but the composition
[…]
Moreover, if $G$ and $X$ are clear from the context, we will usually abbreviate $\mu_{G,X}$ to $\mu$.
Definition 2.3.
A (quasi-)Hamiltonian $G$-variety is a (quasi-)symplectic variety $X$ equipped with a symplectic $G$-action such that the restricted $G$-action to $X_{\operatorname{reg}}$ is Hamiltonian.
Remark 2.4.
The term “quasi-Hamiltonian $G$-spaces” already appears in the literature [AMM98] (in a different context); we emphasize that the notion introduced here is very different from that earlier work.
Given a quasi-Hamiltonian $G$-variety, since $X$ is normal and $\operatorname{codim}(X\setminus X_{\operatorname{reg}})\geq 2$, the moment map $\mu_{G,X_{\operatorname{reg}}}\colon X_{\operatorname{reg}}\rightarrow\mathfrak{g}^{*}$ extends uniquely to a $G$-equivariant morphism $X\rightarrow\mathfrak{g}^{*}$, again denoted by $\mu_{G,X}$. We shall call $\mu_{G,X}$ also the moment map of the quasi-Hamiltonian $G$-variety $X$.
Example 2.5.
Let $Y$ be an irreducible non-singular $G$-variety, $X\coloneqq T^{*}Y$ the cotangent bundle of $Y$. Then $X$ is a non-singular symplectic variety with the canonical symplectic form $d\eta$, where $\eta$ is the Liouville $1$-form on $T^{*}Y$. Moreover, the lifted action of $G$ on $X$ is Hamiltonian, and the moment map is given by
$\mu_{G,X}((y,\alpha),\xi)\coloneqq\langle\alpha,V_{\xi}(y)\rangle,$
where $y\in Y$, $\alpha\in T_{y}^{*}Y$ and $\xi\in\mathfrak{g}$. More generally, singular affine Hamiltonian $G$-varieties can also be constructed by considering the affine closure of cotangent bundles. We refer the reader to [FL25a, FL25b] for examples.
2.2. The shell
Let $X$ be a quasi-Hamiltonian $G$-variety. The shell $N_{X}$ is defined to be the scheme-theoretic fiber of the moment map $\mu\colon X\rightarrow\mathfrak{g}^{*}$ over $0\in\mathfrak{g}^{*}$. We denote by $N_{X,\operatorname{fi}}$ the subset of $N_{X}$ consisting of points $x\in N_{X}\cap X_{\operatorname{reg}}$ such that the isotropy subgroup $G_{x}$ is finite. If $X$ is clear from the context, we will simply denote $N_{X}$ by $N$ and $N_{X,\operatorname{fi}}$ by $N_{\operatorname{fi}}$. The following result is well known for experts, we include a complete proof for the reader’s convenience. See also [KLS06, Lemmas 3.2 and 3.3] and [HSS24b, § 2.3].
Lemma 2.6.
Let $X$ be a Cohen-Macaulay quasi-Hamiltonian $G$-variety such that $N_{\operatorname{fi}}$ is non-empty. Let $n=\dim(X)$ and $d=\dim\mathfrak{g}$.
  1. (2.6.1)
    If $\dim(N\setminus N_{\operatorname{fi}})\leq n-d$, then $N$ is a complete intersection, and hence Cohen-Macaulay. In particular $N$ is non-singular at $x\in X$ if and only if $x\in N_{\operatorname{fi}}$.
  2. (2.6.2)
    If $\dim(N\setminus N_{\operatorname{fi}})\leq n-d-1$, then $N$ is reduced.
  3. (2.6.3)
    If $\dim(N\setminus N_{\operatorname{fi}})\leq n-d-2$, then $N$ is normal.
Proof.
For any point $x\in X$, Condition (H2) implies that for any $v\in T_{x}X$ and any $\xi\in\mathfrak{g}$, we have
$\langle d_{x}\mu(v),\xi\rangle=\omega_{x}(V_{\xi}(x),v).$
Note that $\xi\in\mathfrak{g}_{x}$ if and only if $V_{\xi}(x)=0$, and the latter is equivalent to $\langle d_{x}\mu(v),\xi\rangle=0$ for all $v\in T_{x}X$. It follows that the image of
$d_{x}\mu:T_{x}X\longrightarrow T_{\mu(x)}\mathfrak{g}^{*}\cong\mathfrak{g}^{*}$
is $(\mathfrak{g}/\mathfrak{g}_{x})^{*}\subset\mathfrak{g}^{*}$. Thus, the tangent map $d_{x}\mu$ is surjective if and only if $G_{x}$ is finite. In particular, since $N_{\operatorname{fi}}$ is non-empty, the map $\mu$ is dominant.
For (2.6.1), since $\mu$ is dominant, the shell $N$ is a complete intersection if and only if $N$ has pure dimension ndn-d by [Har77, Theorem 8.21A(c)]. By our assumption and Krull’s Hauptidealsatz, it remains to prove that $\dim(N_{\operatorname{fi}})=n-d$. Note that $X$ is non-singular along $N_{\operatorname{fi}}$ and $\mu$ is smooth along $N_{\operatorname{fi}}$, it follows from [Har77, Proposition 10.4] that $N_{\operatorname{fi}}$ is non-singular of pure dimension ndn-d. So $N$ is Cohen-Macaulay by [Har77, Theorem 8.21A(d)]. Now, it follows from the Jacobian criterion that $N$ is non-singular at $x$ only if $X$ is non-singular at $x$ and hence if and only if $x\in N\cap X_{\operatorname{reg}}$ and $\mu$ is smooth at $x$, which is equivalent to $x\in N_{\operatorname{fi}}$.
For (2.6.2), since the open subset $N_{\operatorname{fi}}$ is dense in $N$ by assumption, the shell $N$ is generically reduced and therefore reduced as it is Cohen-Macaulay.
For (2.6.3), it follows from the assumption that $N$ is Cohen-Macaulay with $\operatorname{codim}(N_{\operatorname{sing}})\geq 2$, so $N$ is normal by Serre’s criterion [Har77, Theorem 8.22A]. ∎
3. Hamiltonian reduction
3.1. Hamiltonian slice theorem
For a scheme $Z$, we denote by $Z^{\operatorname{red}}$ its associated reduced scheme. Let $Y$ be an affine scheme over $\operatorname{Spec}\mathbb{C}$ equipped with a $G$-action. We denote by $Y/\!/G\coloneqq\operatorname{Spec}(\mathcal{O}(Y)^{G})$ the invariant-theoretical quotient with the natural projection $\pi_{Y}\colon Y\rightarrow Y/\!/G$. A subset of $Y$ is $G$-saturated if it is a union of fibers of $\pi_{Y}$. We recall the following basic definition:
Definition 3.1.
Let $X$, $Y$ be affine $G$-varieties. An excellent morphism is a $G$-equivariant morphism $\varphi\colon Y\rightarrow X$ such that the following holds:
  1. (3.1.1)
    The induced morphism $\varphi/\!/G\colon Y/\!/G\rightarrow X/\!/G$ is étale;
  2. (3.1.2)
    The morphism $(\pi_{Y},\varphi)\colon Y\rightarrow Y/\!/G\times_{X/\!/G}X$ is an isomorphism.
In particular, $\varphi\colon Y\rightarrow X$ is also étale, because a base change of an étale map is étale.
[…]
We require the following Hamiltonian slice theorem of Jung [Jun09], which is the symplectic analog of Luna’s slice theorem. A simplified proof of this theorem can be found in [BLLT17, § 6] (see also [HSS20, Theorem 3.15] for the linear case). Although the statement in [BLLT17, Theorem 6.1] differs slightly from the one below, the same argument yields our version. For completeness, we include a detailed proof in Appendix § A with some minor adaptations to the details.
Theorem 3.2.
Let $X$ be an affine quasi-Hamiltonian $G$-variety equipped with a symplectic form $\omega$ on $X_{\operatorname{reg}}$. Let $O_{x}=Gx$ be a closed orbit with $x\in N_{X}\cap X_{\operatorname{reg}}$. Then there exists a non-singular locally closed affine $G_{x}$-stable subset $U\subset X_{\operatorname{reg}}$ containing $x$ such that $\omega|_{U}$ is symplectic and the morphism
$\psi\colon G\times_{G_{x}}N^{\operatorname{red}}_{U}\longrightarrow N^{\operatorname{red}}_{X},\quad(g,u)\longmapsto gu$
is excellent, where $N_{U}$ is the shell of the Hamiltonian $G_{x}$-variety $U$.
The locally closed subset $U\subset X_{\operatorname{reg}}$ with the properties of the theorem will be called a Hamiltonian slice at $x$.
3.2. Isotropy-type stratification
Let $Y$ be an affine $G$-variety (not necessarily irreducible). For $r\in\mathbb{Z}_{\geq 0}$, we denote by $Y_{(r)}$ the subset of $Y$ consisting of points with isotropy groups of dimension $r$, which is a locally closed subset by the upper semi-continuity of $\dim(G_{y})$, $y\in Y$.
We denote by $(G_{y})$ the conjugacy class of $G_{y}$, $y\in Y$, which we call an isotropy class of $Y$, and an isotropy class is called closed if the orbit $O_{y}=Gy$ is closed in $Y$. Then a closed isotropy class is reductive by Matsushima’s theorem, and there are only finitely many closed isotropy classes of $Y$ by Luna’s slice theorem [Lun73]. For a closed isotropy class $(H)$ of $G$, we denote by $(Y/\!/G)_{(H)}$ the points in Y//GY/\!/G corresponding to closed orbits with isotropy class in $(H)$. This is a constructible subset of Y//GY/\!/G by Luna’s slice theorem. Let $Y^{(H)}$ be the inverse image of $(Y/\!/G)_{(H)}$ in $Y$.
Assume in addition that Y//GY/\!/G is irreducible, then there is a unique closed isotropy class $(H)$, called the principal isotropy class, such that $(Y/\!/G)_{\operatorname{pr}}\coloneqq(Y/\!/G)_{(H)}$ is dense in Y//GY/\!/G and the closed orbits $O_{y}$ with $G_{y}\in(H)$ are called principal orbits. Then $(Y/\!/G)_{\operatorname{pr}}\subset Y/\!/G$ is called the principal stratum and is open in Y//GY/\!/G by Luna’s slice theorem. Let $Y_{\operatorname{pr}}\coloneqq\pi_{Y}^{-1}((Y/\!/G)_{\operatorname{pr}})$. We say that $Y$ has finite principal isotropy groups (FPIG) if its principal isotropy groups are finite.
3.3. Hamiltonian reduction
Definition 3.3.
The Hamiltonian reduction of an affine Hamiltonian $G$-variety $X$ is defined as
$X/\!\!/\!\!/G\coloneqq N_{X}/\!/G.$
Now we apply Theorem 3.2 to study the isotropy-type stratifications of Hamiltonian reductions.
Lemma 3.4.
Let $X$ be an affine Hamiltonian $G$-variety. Let $O_{x}=Gx\subset N_{X}$ be a closed orbit with $x\in N_{X}\cap X_{\operatorname{reg}}$ . Let $x\in U$ be a Hamiltonian slice. Then there exists an open subset $V$ of $N_{X}^{\operatorname{red}}/\!/G$ containing $O_{x}]$ such that
$\dim\left(V\cap(N_{X}^{\operatorname{red}}/\!/G)_{(G_{x})}\right)=\dim(N_{U}^{\operatorname{red}}/\!/G_{x})_{(G_{x})}.$
Proof.
By Theorem 3.2 and [BKS25, Proposition 5.3(2)], there exists an étale morphism
$\psi/\!/G\colon N_{U}^{\operatorname{red}}/\!/G_{x}\cong(G\times_{G_{x}}N_{U}^{\operatorname{red}})/\!/G\longrightarrow N_{X}^{\operatorname{red}}/\!/G.$
Denote by $V$ the image of $\psi/\!/G$. Then $V$ is open and contains $O_{x}]$. Moreover, by the second condition of excellent morphisms, we have
$(\psi/\!/G)^{-1}\left((N_{X}^{\operatorname{red}}/\!/G)_{(G_{x})}\right)=\left((G\times_{G_{x}}N_{U}^{\operatorname{red}})/\!/G\right)_{(G_{x})}=\left(N_{U}^{\operatorname{red}}/\!/G_{x}\right)_{(G_{x})},$
from which the required equality follows. ∎
Lemma 3.5.
Let $X$ be an affine Hamiltonian $G$-variety. Let $O_{x}=Gx\subset N_{X}$ be a closed orbit with $x\in N_{X}\cap X_{\operatorname{reg}}$ . Let $x\in U$ be a Hamiltonian slice. Then we have
$\dim(N_{U}^{\operatorname{red}}/\!/G_{x})_{(G_{x})}=\dim(T_{x}X)^{G_{x}}-2(\dim{\rm Norm}_{G}(G_{x})-\dim G_{x}),$
where ${\rm Norm}_{G}(G_{x})$ is the normalizer of $G_{x}$ in $G$.
Proof.
Note that $U^{G_{x}}$ is contained in $N_{U}$ and the induced morphism
$U^{G_{x}}\longrightarrow(N_{U}^{\operatorname{red}}/\!/G_{x})_{(G_{x})}$
is bijective, so one gets
$\dim(N_{U}^{\operatorname{red}}/\!/G_{x})_{(G_{x})}=\dim\left(N_{U}^{\operatorname{red}}\right)^{G_{x}}=\dim U^{G_{x}}.$
On the other hand, by Corollary A.5, we have
$\dim(U^{G_{x}})=\dim(T_{x}U)^{G_{x}}=\dim(T_{x}X)^{G_{x}}-\dim(\mathfrak{g}/\mathfrak{g}_{x})^{G_{x}}-\dim((\mathfrak{g}/\mathfrak{g}_{x})^{*})^{G_{x}}.$
In particular, since
$\dim(\mathfrak{g}/\mathfrak{g}_{x})^{G_{x}}=\dim((\mathfrak{g}/\mathfrak{g}_{x})^{*})^{G_{x}}=\dim{\rm Norm}_{G}(G_{x})-\dim(G_{x}),$
the lemma follows. ∎
Lemma 3.6.
Let $X$ be a non-singular affine Hamiltonian $G$-variety, equipped with a symplectic form $\omega$. Assume that $N$ is reduced and has FPIG such that N//GN/\!/G is irreducible. Then $(N/\!/G)_{\operatorname{pr}}\subset(N/\!/G)_{\operatorname{reg}}$ and there exists a natural symplectic form on $(N/\!/G)_{\operatorname{pr}}$.
Proof.
Let $O_{x}=Gx\subset N$ be a principal orbit and let $x\in U\subset X$ be a Hamiltonian slice. Since $G_{x}$ is finite, we have $N_{U}=U$. Consider the excellent $G_{x}$-equivariant morphism $U\rightarrow T_{x}U$ (see Corollary A.5). Since $(G_{x})$ is the principal isotropy class, the set of points in $U$ with stabilizer $G_{x}$ is open and dense in $U$, in particular, $U^{G_{x}}$ is open and dense. This implies that the morphisms appearing in the following commutative diagram are dominant:
$U^{G_{x}}$$(T_{x}U)^{G_{x}}$$U/\!/G_{x}$$(T_{x}U)/\!/G_{x}.$}
So $\dim(T_{x}U)^{G_{x}}=\dim T_{x}U$ and hence $G_{x}$ acts trivially on $T_{x}U$ and $U$. It follows that $U/\!/G_{x}=U$ is non-singular and carries a natural symplectic form. Finally, we observe that there exists a natural symplectic linear isomorphism
$T_{x}U\longrightarrow(T_{x}O_{x})^{\perp_{\omega_{x}}}/T_{x}O_{x},$
so, the symplectic form is independent of the choice of the Hamiltonian slice $U$. ∎
The following criterion allows us to determine when a Hamiltonian reduction is a symplectic variety.
Proposition 3.7.
Let $X$ be a non-singular affine Hamiltonian $G$-variety. Assume that
  1. (3.7.1)
    $N_{\operatorname{fi}}$ is non-empty,
  2. (3.7.2)
    $\dim(N\setminus N_{\operatorname{fi}})\leq\dim(X)-\dim(G)-2$,
  3. (3.7.3)
    $N^{\operatorname{red}}$ has FPIG,
  4. (3.7.4)
    $N^{\operatorname{red}}/\!/G$ is irreducible,
  5. (3.7.5)
    for any $x\in N^{\operatorname{red}}\setminus N^{\operatorname{red}}_{\operatorname{pr}}$ with closed $G$-orbit, we have
    $\operatorname{codim}(T_{x}X)^{G_{x}}\geq 2(\dim G-\dim{\rm Norm}_{G}(G_{x})+\dim G_{x})+4.$
Then X///GX/\!\!/\!\!/G is a symplectic variety with terminal singularities.
Proof.
By Lemma 2.6, the shell $N$ is normal, so $N=N^{\operatorname{red}}$ and consequently N//GN/\!/G is also normal. In particular, since $N$ has FPIG by assumption, the open subset $N_{\operatorname{pr}}$ consists precisely of the closed principal orbits by Luna’s slice theorem. This implies
$\dim(X/\!\!/\!\!/G)=\dim(N/\!/G)=\dim N-\dim G=\dim X-2\dim G.$
Now, let $(H)$ be a closed non-principal isotropy class. By Lemmas 3.4, 3.5 and the assumption, we have
$\dim(N/\!/G)_{(H)}\leq\dim X-2\dim G-4\leq\dim(N/\!/G)-4.$
This implies that $(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codimension at least four, and the canonical symplectic form on $(N/\!/G)_{\operatorname{pr}}$ obtained in Lemma 3.6 then extends to a $2$-form on $(N/\!/G)_{\operatorname{reg}}$, which is again symplectic because its degenerate locus has codimension one if it is non-empty. Finally, since $(N/\!/G)\setminus(N/\!/G)_{\operatorname{reg}}\subset(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codimension at least four, it follows that X///G=N//GX/\!\!/\!\!/G=N/\!/G is a symplectic variety by Flenner’s theorem [Fle88] and with terminal singularities by [Nam01]. ∎
Remark 3.8.
A modified version of Proposition 3.7 also holds for Cohen–Macaulay affine quasi-Hamiltonian $G$-varieties. The main difficulty in the singular setting is that the Hamiltonian slice theorem does not hold at singular points, and we have no control on the dimension of $(N/\!/G)_{(H)}$ at closed orbits lying in the locus $N\cap X_{\operatorname{sing}}$. However, if the image $\pi_{N}(N\cap X_{\operatorname{sing}})$ has codimension at least four in N//GN/\!/G (for example, when $\dim X_{\operatorname{sing}}\leq\dim X-2\dim G-4$), then Proposition 3.7 remains valid when we restrict it to points $x\in(N^{\operatorname{red}}\cap X_{\operatorname{reg}})\setminus N_{\operatorname{pr}}^{\operatorname{red}}$. We leave the verification of these details to the interested reader.
4. Hamiltonian reduction of cotangent bundles
We now apply Proposition 3.7 to study the Hamiltonian reduction of the cotangent bundle of a smooth affine variety with a $2$-large $G$-action, thus proving Theorems 1.1 and 1.3.
Throughout this section, we fix the following notation. Given an irreducible non-singular affine $G$-variety $Y$, we denote by $\mu\colon T^{*}Y\rightarrow\mathfrak{g}^{*}$ the canonical moment map. The shell of $\mu$ is denoted by $N$. The GIT quotient of $Y$ under $G$ is written as $Z\coloneqq Y/\!/G$, with the quotient map $\pi_{Y}\colon Y\rightarrow Z$. The Hamiltonian reduction of $T^{*}Y$ is denoted as $M\coloneqq T^{*}Y/\!\!/\!\!/G=N/\!/G$ with the corresponding quotient $\pi_{N}\colon N\rightarrow M$.
4.1. Large actions
We start with some basic definitions.
Definition 4.1.
Let $Y$ be an irreducible affine $G$-variety and $k\in\mathbb{Z}_{\geq 0}$. Then $Y$ is called
  1. (4.1.1)
    $k$-principal if $\operatorname{codim}(Y\setminus Y_{\operatorname{pr}})\geq k$, and
  2. (4.1.2)
    $k$-modular if $\operatorname{codim}(Y_{(r)})\geq r+k$ for $1\leq r\leq\dim(G)$, and
  3. (4.1.3)
    $k$-large if $Y$ is $k$-principal, $k$-modular and has FPIG.
The following result is a non-linear generalization of [HSS20, Proposition 3.2].
Proposition 4.2.
Let $Y$ be an irreducible non-singular affine $G$-variety. Let $\dim(Y)=n$, $\dim(G)=d$ and $k\in\mathbb{Z}_{\geq 0}$.
  1. (4.2.1)
    If $Y$ is $k$-modular, then $N_{\operatorname{fi}}$ is non-empty and $\dim(N\setminus N_{\operatorname{fi}})\leq 2n-d-k$. In particular, the shell $N$ is a Cohen-Macaulay complete intersection.
  2. (4.2.2)
    If $Y$ is $1$-modular, then $N$ is irreducible and reduced.
  3. (4.2.3)
    If $Y$ is $2$-modular, then $N$ is irreducible and normal.
Proof.
For points $y\in Y$ and $w\in T_{y}^{*}Y$, we have $G_{w}\subset G_{y}$. In particular, if $Y$ is $k$-modular, then $Y_{(0)}$ is non-empty and consequently $N_{\operatorname{fi}}$ is non-empty. On the other hand, note that the restricted map
$\mu|_{T_{y}^{*}Y}\colon T_{y}^{*}Y\longrightarrow\mathfrak{g}^{*}$
is exactly given by the cotangent map at $y$ of the orbit map $\operatorname{ev}_{y}\colon G\rightarrow Gy\eqqcolon O_{y}\subset Y$, $g\mapsto gy$; that is, the dual of the tangent map
$d_{e}\operatorname{ev}_{y}\colon T_{e}G=\mathfrak{g}\longrightarrow T_{y}O_{y}\subset T_{y}Y.$
In particular, the intersection $N\cap T_{y}^{*}Y$ is a linear subspace of $T_{y}^{*}Y$ with codimension $\dim O_{y}=d-\dim(G_{y})$. Let $p\colon T^{*}Y\rightarrow Y$ be the natural projection. Set $N^{r}=p^{-1}(Y_{(r)})\cap N$. Then, for $1\leq r\leq\dim G$, we have
(4.2.1) $\dim N^{r}\leq\dim Y_{(r)}+(n-d+r)\leq 2n-d-k.$
In particular, since $N_{\operatorname{fi}}$ contains $N^{0}$, one obtains
$\dim(N\setminus N_{\operatorname{fi}})\leq\max_{1\leq r\leq d}\dim N^{r}\leq 2n-d-k.$
Now, by Lemma 2.6, it remains to prove the irreducibility of $N$ for $k\geq 1$. Note that the fibers of the projection
$N^{0}\longrightarrow Y_{(0)}$
is a codimension $d$ linear subspace of $T^{*}_{y}Y$. In particular, since $Y_{(0)}$ is irreducible, the variety $N^{0}$ is also irreducible. Now we observe
$\dim(N\setminus N^{0})\leq\max_{1\leq r\leq d}\dim N^{r}\leq 2n-d-k\leq 2n-d-1<\dim N.$
Hence $N$ is irreducible as it has pure dimension 2nd2n-d. ∎
Lemma 4.3 (cf. [HSS20, § 4.1]).
Let $Y$ be an irreducible non-singular $2$-large affine $G$-variety.
  1. (4.3.1)
    Then $N$ is an irreducible normal variety such that $G_{x}=G_{y}$ for any $x\in N$ with $p(x)\in Y_{\operatorname{pr}}$, where $p\colon T^{*}Y\rightarrow Y$ is the natural projection. In particular, $N$ has FPIG and satisfies
    $p^{-1}(Y_{\operatorname{pr}})\cap N\subset N_{\operatorname{pr}}.$
  2. (4.3.2)
    For any $x\in N\setminus N_{\operatorname{pr}}$ with closed $G$-orbit, we have
    $\operatorname{codim}(T_{x}(T^{*}Y))^{G_{x}}\geq 2(\dim G-\dim{\rm Norm}_{G}(G_{x})+\dim G_{x})+4.$
Proof.
Firstly, we note that $N$ is an irreducible normal variety by Proposition 4.2. Moreover, since $Y$ has FPIG, to prove (4.3.1), it suffices to show that any point $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ has a closed $G$-orbit such that $G_{x}=G_{y}$.
Let $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ and $y\coloneqq p(x)\in Y_{\operatorname{pr}}$. Moreover, since $Y$ has FPIG, the open subset $Y_{\operatorname{pr}}$ consists of closed principal orbits, so $O_{y}=Gy$ is closed in $Y$ and $G_{x}\subset G_{y}$ is a finite subgroup. Moreover, the orbit $O_{x}=Gx$ is contained in the closed subset
$p^{-1}(O_{y})\cong G\times_{G_{y}}T_{y}^{*}Y$
of $X$ such that $O_{x}\cong G\times_{G_{y}}O^{\prime}_{x}$, where $G_{y}$ acts on $T^{*}_{y}Y$ by the isotropy action and the orbit $O^{\prime}_{x}=G_{y}x\subset T_{y}^{*}Y$ is a finite set. So $O_{x}$ is closed in $p^{-1}(O_{y})$ and is therefore closed in $T^{*}Y$.
Since $G_{y}$ is finite and $y\in Y_{\operatorname{pr}}$, it follows from Luna’s slice theorem that there exists an étale slice $S\subset Y$ at $y$ such that $G_{y}$ acts on $S$ trivially (cf. Proof of Lemma 3.6 and [BKS25, Theorem 5.3]). In particular, $G_{y}$ acts on $T_{y}S$ trivially. In particular, as $N\cap T_{y}^{*}Y\cong T_{y}^{*}S$ as $G_{y}$-modules, one derives $G_{x}=G_{y}$, hence the statement (4.3.1) follows.
For (4.3.2), consider any $x\in N\setminus N_{{\rm pr}}$ with closed orbit so that $y=p(x)\not\in Y_{\operatorname{pr}}$. Following the arguments in [HSS20, Lemmas 4.1, 4.2 and 4.3], we have
$\dim Y^{G_{x}}\leq\dim Y-\dim G+\dim{\rm Norm}_{G}(G_{x})-\dim G_{x}-2,$
which then implies
$\operatorname{codim}(T_{x}(T^{*}Y))^{G_{x}}$ $=2(\dim Y-\dim Y^{G_{x}})$
$\geq 2(\dim G-\dim{\rm Norm}_{G}(G_{x})+\dim G_{x})+4.$
as desired. ∎
Proof of Theorem 1.1.
The statement (1.1.1) follows from Proposition 3.7, Proposition 4.2 and Lemma 4.3. For simplicity, we set $N^{\diamond}=p^{-1}(Y_{\operatorname{pr}})\cap N$ $(\subset N_{\operatorname{pr}})$ and denote $\pi_{N}(N^{\diamond})$ by $M^{\diamond}$.
For (1.1.2), since $Y$ and $N$ have FPIG, by Luna’s slice theorem, the principal strata $Z_{\operatorname{pr}}$ and $M_{\operatorname{pr}}$ are non-singular such that the induced morphisms $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ and $N_{\operatorname{pr}}\rightarrow M_{\operatorname{pr}}$ are $G$-fiber bundles. In particular $M^{\diamond}$ is a dense open subset of $M$. Then the cotangent map $d^{*}\pi_{Y}$ of $\pi_{Y}$ induces an injective morphism $\pi_{Y}^{*}(T^{*}Z_{\operatorname{pr}})\to T^{*}Y_{\operatorname{pr}}$ whose image is exactly $N^{\diamond}$. This gives the following commutative diagram with isomorphic horizontal morphisms:
$\pi^{*}(T^{*}Z_{\operatorname{pr}})$$N^{\diamond}$$T^{*}Z_{\operatorname{pr}}$$M^{\diamond}.$$\scriptstyle{d^{*}\pi_{Y}}$$\scriptstyle{\pi_{N}}$$\scriptstyle{\iota}$
More precisely, the map $\iota$ can be explicitly described as follows: for any $y\in Y_{{\rm pr}}$ with $z=\pi_{Y}(y)\in Z_{\operatorname{pr}}$ and $\eta\in T^{*}_{z}Z_{\operatorname{pr}}$, we define $\tilde{\eta}\in T_{y}^{*}Y$ by taking the composition
[…]
Note that $T_{y}O_{y}$ is contained in the kernel of $d_{y}\pi_{Y}$, we have $\tilde{\eta}(T_{y}O_{y})=0$, which shows that $\tilde{\eta}\in N^{\diamond}$. Then one easily checks that the following morphism is a well-defined isomorphism (i.e. independent of the choice of $y\in\pi_{Y}^{-1}(z))$:
[…]
Since $Y$ is $2$-principal, the closed subset $Y\setminus Y_{\operatorname{pr}}$ has codimension at least two in $Y$, so that $Z_{\operatorname{reg}}\setminus Z_{\operatorname{pr}}$ has codimension at least two in $Z_{\operatorname{reg}}$. This implies that the restriction $\mathcal{O}(T^{*}Z_{\operatorname{reg}})\rightarrow\mathcal{O}(T^{*}Z_{\operatorname{pr}})$ is an isomorphism of $\mathbb{C}$-algebras. So it remains to prove that $\operatorname{codim}M\setminus M^{\diamond}\geq 2$. Let $N^{r}\coloneqq p^{-1}(Y_{(r)})\cap N$ for any $0\leq r\leq\dim G$. Then, for $1\leq r\leq\dim G$, by (4.2.1), we have
$\dim N^{r}\leq 2\dim Y-\dim G-2.$
On the other hand, since $Y$ is $2$-principal, we have
$\dim N^{0}\setminus N^{\diamond}=\dim(Y_{(0)}\setminus Y_{\operatorname{pr}})+\dim Y-\dim G\leq 2\dim Y-\dim G-2.$
This yields
$\dim N\setminus N^{\diamond}\leq 2\dim Y-\dim G-2=\dim N-2,$
which implies $\dim(M\setminus M^{\diamond})\leq\dim M-2$, and hence (1.1.2) follows.
For (1.1.3), firstly note that if $Z$ is non-singular, then $T^{*}Z$ is an irreducible non-singular affine variety such that $M=T^{*}Z$, so $M$ is also non-singular.
For the converse, we assume that $Z$ is singular. By (1.1.2), there exists a natural projection $p_{Z}\colon M\rightarrow Z$ such that the restriction $p_{Z}|_{T^{*}Z_{\operatorname{pr}}}\colon T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ identifies to the natural projection. Let $\mathbf{Z}$ be the zero section of $p_{Z}$; that is, the closure of the zero section $\mathbf{Z}_{\operatorname{pr}}$ of $T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ in $M$ or equivalently $M^{\mathbb{C}^{*}}$, where the $\mathbb{C}^{*}$-action on $M$ is induced by the natural scaling on the fibers of $T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$. Then the restriction $\iota=p_{Z}|_{\mathbf{Z}}\colon\mathbf{Z}\longrightarrow Z$ is an isomorphism.
Let $z$ be a singular point of $Z$ and let $\mathbf{z}\in\mathbf{Z}$ be the point such that $\iota(\mathbf{z})=z$. Let $F$ be an irreducible component of the fiber of $p_{Z}$ passing through $\mathbf{z}$. Then $\dim F\geq\dim Z$ by the upper semi-continuity. Let $T_{z}Z$, $T_{\mathbf{z}}M$ and $T_{\mathbf{z}}F$ be the Zariski tangent spaces. Then we have the following complex of $\mathbb{C}$-linear maps given by the tangent maps:
$T_{\mathbf{z}}F\longrightarrow T_{\mathbf{z}}M\longrightarrow T_{z}Z,$
where the first map is injective. Moreover, since $\iota\colon\mathbf{Z}\rightarrow Z$ is an isomorphism, the following composition
$T_{\mathbf{z}}\mathbf{Z}\longrightarrow T_{\mathbf{z}}M\longrightarrow T_{z}Z$
is an isomorphism, so the last map is surjective. On the other hand, since $z$ is a singular point of $Z$, we have $\dim T_{z}Z>\dim Z$, which implies
$\dim T_{\mathbf{z}}M\geq\dim T_{\mathbf{z}}F+\dim T_{z}Z>\dim F+\dim Z\geq 2\dim Z=\dim M.$
Hence $M$ is singular at $\mathbf{z}$ by Jacobian’s criterion. ∎
Remark 4.4.
For an irreducible non-singular affine $G$-variety $Y$ that is $2$-large, we have $Z_{\operatorname{pr}}=Z_{\operatorname{reg}}$ by [Sch95, Theorem 9.12]. So $Z$ is non-singular if and only if $Z=Z_{\operatorname{pr}}$. Then Luna’s slice theorem implies that $Y\rightarrow Z$ is actually a $G$-fiber bundle with fiber $G/H$ (see [BKS25, Proposition 6.4]), where $(H)$ is the principal isotropy class.
Example 4.5.
Let $Y$ be an irreducible non-singular affine $G$-variety with FIPG. Assume that the identity component $G^{0}$ is a torus. By [Sch95, Proposition 10.1], $Y$ is $k$-large if and only if $Y$ is $k$-principal. If $G$ acts on $Y$ freely outside a locus of codimension at least two, then $Y$ is $2$-large. Theorem 1.1 implies that $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities that is the affine closure of $T^{*}(Y/\!/G)_{\operatorname{pr}}$. One particular example is that when $G$ is a finite group that acts freely on $Y^{\diamond}\subset Y$ with $\operatorname{codim}Y\setminus Y^{\diamond}\geq 2$, then $T^{*}Y/G$ is the affine closure of $T^{*}(Y^{\diamond}/G)$, as previously remarked in [FL25a, Example 2.8].
4.2. Factoriality
Recall from [KKV89] that for a $G$-variety $Y$, ${\rm Pic}_{G}(Y)$ denotes the set of isomorphism classes of line bundles on $Y$ with a $G$-linearization. Forgetting the $G$-linearization, we have the canonical homomorphism $\nu\colon{\rm Pic}_{G}(Y)\to{\rm Pic}(Y)$. The kernel ${\rm ker}(\nu)$ consists of $G$-linearizations of the trivial line bundle up to isomorphisms, which is denoted by $H^{1}_{\rm alg}(G,\mathcal{O}(Y)^{\times})$, where $\mathcal{O}(Y)^{\times}$ is the group of invertible regular functions on $Y$.
For an irreducible non-singular affine $G$-variety $Y$, we remark that it follows from the standard properties of affine GIT that the natural projections $Y\rightarrow Z$ and $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ are both quotients by $G$ in the sense of [KKV89, 4.1].
Lemma 4.6.
Let $Y$ be an irreducible $2$-large non-singular affine $G$-variety.
  1. (4.6.1)
    If both $\operatorname{Pic}(Y)$ and $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ are finite, then $\operatorname{Cl}(Z)$ is finite and hence $Z$ is $\mathbb{Q}$-factorial.
  2. (4.6.2)
    If both $\operatorname{Pic}(Y)$ and $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ are trivial, then $\operatorname{Cl}(Z)$ is trivial and hence $Z$ is factorial.
Proof.
Since $Y$ is $2$-principal, it follows that $\operatorname{codim}(Y\setminus Y_{\operatorname{pr}})\geq 2$ for the $G$-stable open subset $Y_{\operatorname{pr}}$ of $Y$. In particular, one gets
$\operatorname{Pic}(Y_{\operatorname{pr}})\cong\operatorname{Pic}(Y)\quad\text{and}\quad\mathcal{O}(Y)^{\times}=\mathcal{O}(Y_{\operatorname{pr}})^{\times}.$
Moreover, since $Z\setminus Z_{\operatorname{pr}}$ has codimension at least two in $Z$, $Z_{\operatorname{pr}}\subset Z_{\operatorname{reg}}$ and $Z$ is normal, we have $\operatorname{Cl}(Z)\cong\operatorname{Pic}(Z_{\operatorname{pr}})$. By [KKV89, Lemma 2.2], there exists an exact sequence
$1\to H^{1}_{\rm alg}(G,\mathcal{O}(Y_{\operatorname{pr}})^{\times})\to{\rm Pic}_{G}(Y_{\operatorname{pr}})\to{\rm Pic}(Y_{\operatorname{pr}}).$
Then applying [KKV89, Proposition 4.2] to the quotient $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ shows that $\operatorname{Pic}(Z_{\operatorname{pr}})$ is a subgroup of $\operatorname{Pic}_{G}(Y_{\operatorname{pr}})$, so the result follows. ∎
Proof of Theorem 1.3.
Firstly note that we have natural isomorphisms of groups
$\operatorname{Cl}(T^{*}Y/\!\!/\!\!/G)\cong\operatorname{Pic}(T^{*}Z_{\operatorname{pr}})\cong\operatorname{Pic}(Z_{\operatorname{pr}})\cong\operatorname{Cl}(Z),$
where the first isomorphism follows from Theorem 1.1. By Lemma 4.6, $\operatorname{Cl}(Z)$ is trivial (resp. finite) if both $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ and $\operatorname{Pic}(Y)$ are trivial (resp. finite). By assumption, either $E(Y)\coloneqq\mathcal{O}(Y)^{\times}/\mathbb{C}^{\times}$ is trivial or $G$ is connected, so $H^{1}(G/G^{0},E(Y))$ is trivial. In particular, the following exact sequence from [KKV89, Proposition 2.3]
$\mathfrak{X}(G)\to H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})\to H^{1}(G/G^{0},E(Y)).$
implies that $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ is trivial (resp. finite) if the character group $\mathfrak{X}(G)$ of $G$ is trivial (resp. finite) and hence if $G_{\operatorname{ab}}$ is trivial (resp. finite). ∎
Proof of Theorem 1.4.
Recall that the singular locus of terminal singularities has codimension at least three and it follows from [Fu03, Corollary 1.3] that a locally $\mathbb{Q}$-factorial singular symplectic variety does not admit any symplectic resolution if its singular locus is not of pure codimension two.
Since $G$ is semisimple by assumption, it follows that $G$ is connected and $G_{\operatorname{ab}}$ is trivial, therefore $T^{*}Y/\!\!/\!\!/G$ is a $\mathbb{Q}$-factorial symplectic variety with terminal singularities by Theorems 1.1 and 1.3. In particular, $T^{*}Y/\!\!/\!\!/G$ admits a symplectic resolution if and only if $T^{*}Y/\!\!/\!\!/G$ itself is non-singular, and hence if and only if Y//GY/\!/G is non-singular by Theorem 1.1. ∎
Under some additional assumptions on $Y$, Theorem 1.3 can be slightly improved.
Proposition 4.7.
Let $Y$ be an irreducible non-singular affine $G$-variety that is $2$-large. Assume that $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$, $Y^{G}\not$$\emptyset$ and $\operatorname{Pic}(Y)=0$.
  1. (4.7.1)
    The quotient $M$ is factorial (resp. $\mathbb{Q}$-factorial) if and only if $\operatorname{Cl}(Z)$ is trivial (resp. a torsion group).
  2. (4.7.2)
    The divisor class group $\operatorname{Cl}(Z)$ is isomorphic to the kernel of $\mathfrak{X}(G)\rightarrow\mathfrak{X}(H)$, where $(H)$ is the principal isotropy class.
Proof.
Applying [KKV89, Corollary 5.3] to the quotient $Y\rightarrow Z$ yields $\operatorname{Pic}(Z)=0$. Let $p_{Z}\colon M\rightarrow Z$ be the natural projection and let $\mathbf{Z}$ be the zero section of $p_{Z}$ with the isomorphism $\iota=p_{Z}|_{\mathbf{Z}}\colon\mathbf{Z}\longrightarrow Z$. Consider the following commutative diagram:
$0=\operatorname{Pic}(\mathbf{Z})$$\operatorname{Pic}(M)$$\operatorname{Cl}(T^{*}Z_{\operatorname{pr}})=\operatorname{Cl}(M)$$0=\operatorname{Pic}(Z)$$\operatorname{Cl}(Z_{\operatorname{pr}})=\operatorname{Cl}(Z).$$\scriptstyle{p_{Z}^{*}}\scriptstyle{\cong}$$\scriptstyle{\iota^{*}}$$\scriptstyle{p_{Z}^{*}}$
Then one easily derives that for a Weil divisor $D$ on $Z$, the pull-back $p_{Z}^{*}D$ is Cartier (resp. $\mathbb{Q}$-Cartier) if and only if $D$ is Cartier (resp. $\mathbb{Q}$-Cartier) and if and only if $D$ is trivial (resp. a torsion) in $\operatorname{Cl}(Z)$, and hence (4.7.1) follows.
Finally, as $\operatorname{Cl}(Z)=\operatorname{Pic}(Z_{\operatorname{pr}})$, we may apply [KKV89, Proposition 5.1 and Remark 4.5] to the quotient $Y_{\operatorname{pr}}\to Z_{\operatorname{pr}}$ to obtain the exact sequence
$1\longrightarrow\operatorname{Pic}(Z_{\operatorname{pr}})\longrightarrow\mathfrak{X}(G)\longrightarrow\mathfrak{X}(H),$
which then gives the desired statement (4.7.2). ∎
As an application, we prove the following partial converse to Theorem 1.3 (for related results in the linear case, see [BS19, Theorem 1.2 and Corollary 2.8]).
Corollary 4.8.
Under the notation and assumptions of Proposition 4.7, the following hold:
  1. (4.8.1)
    The quotient $M$ is $\mathbb{Q}$-factorial if and only if $G_{\operatorname{ab}}$ is finite.
  2. (4.8.2)
    If $H$ is trivial, then $M$ is factorial if and only if $G$ is perfect.
Proof.
By Theorem 1.3 and Proposition 4.7, we only need to prove that if $M$ is $\mathbb{Q}$-factorial, then $G_{\operatorname{ab}}$ is finite. Assume to the contrary that $G_{\operatorname{ab}}$ is not finite. Then we can choose a surjective character $\chi\colon G\rightarrow\mathbb{C}^{\times}$. In particular, $\chi^{n}\not=1$ for all $n\geq 1$. Nevertheless, since the principal isotropy gruop $H$ is finite, there exists $m\in\mathbb{Z}_{>0}$ such that $\chi^{m}|_{H}=1$, which then implies that $\operatorname{Cl}(Z)$ is not torsion and hence $M$ is not $\mathbb{Q}$-factorial by Proposition 4.7, which is absurd. ∎
The following result generalizes [BS19, Corollary 1.3].
Corollary 4.9.
Under the notation and assumptions of Proposition 4.7, if $G_{\operatorname{ab}}$ is finite and $G$ acts non-trivially on $Y$, then $M$ does not admit a symplectic resolution.
Proof.
By Theorem 1.1 and Theorem 1.3, it remains to show that $Z$ is singular. Indeed, by [Sch95, Theorem 9.12], we have $Z_{\operatorname{reg}}=Z_{\operatorname{pr}}$. In particular, if $Z$ is non-singular, then $Z_{\operatorname{pr}}=Z$ and $Y_{\operatorname{pr}}=Y$, and it follows that $G$ is a finite group acting trivially on $Y$ as $Y^{G}\not$$\emptyset$, which is a contradiction. ∎
4.3. Examples
We gather some examples to illustrate the sharpness of Theorems 1.1 and 1.3.
Example 4.10 ([BS19, Example 3.5]).
Let $V=(\mathfrak{sl}_{2})^{2}$, as a representation of $G=\operatorname{PGL}_{2}$, which is $1$-large, but not $2$-large. The Hamiltonian reduction $T^{*}V/\!\!/\!\!/G$ is identified with the locus of square zero matrices in $\mathfrak{sp}_{4}$, and in particular is a symplectic variety. Nevertheless, its singular locus is the codimension two locus consisting of rank one matrices in $\mathfrak{sp}_{4}$ and therefore it is not terminal, so the part (1.1.1) of Theorem 1.1 is optimal.
Example 4.11.
For $0\leq k\leq n$, let $G=\mathbb{C}^{\times}$ act on $V_{k}=\mathbb{C}^{n}$ by
$\lambda\cdot(x_{1},\dots,x_{n})=(\lambda x_{1},\dots,\lambda x_{k},\lambda^{-1}x_{k+1},\cdots,\lambda^{-1}x_{n}).$
It is easy to see that for any $k$, the Hamiltonian reduction $T^{*}V_{k}/\!\!/\!\!/G$ is isomorphic to the minimal nilpotent orbit closure $\overline{\mathbf{O}}_{{\rm min}}$ in $\mathfrak{sl}_{n}$. For $z\in V_{k}$, the $G$-orbit GzGz is closed in $V_{k}$ if and only if $z=0$ or $z\in V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, where $F_{k}=\{x_{1}=\cdots=x_{k}=0\}$ and $F^{\prime}_{n-k}=\{x_{k+1}=\cdots=x_{n}=0\}$. It follows that $(V_{k})_{\operatorname{pr}}=V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, so $V_{k}$ is 2-large if and only if $2\leq k\leq n-2$. In this case, Theorem 1.1 implies that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of $T^{*}((V_{k}/\!/G)_{\operatorname{pr}})$. On the other hand, $(V_{k}/\!/G)_{\operatorname{pr}}$ is just the punctured affine cone of the Segre variety $\mathbb{P}^{k-1}\times\mathbb{P}^{n-k-1}\subset\mathbb{P}^{k(n-k)-1}$. This shows that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of the cotangent bundle of the punctured affine cone of a Segre variety, which is a special case of [FL25a, Theorem 1.4] but with a new proof. For $k=0$, $V_{0}/\!/G$ is just one point. For $k=1$, $V_{1}$ is $1$-large with $V_{1}/\!/G\cong\mathbb{C}^{n-1}$ and $(V_{1}/\!/G)_{\operatorname{pr}}\cong\mathbb{C}^{n-1}\setminus\{0\}$. It follows that the affine closure of $T^{*}(V_{1}/\!/G)_{\operatorname{pr}}$ is $T^{*}\mathbb{C}^{n-1}$ which is different from $\overline{\mathbf{O}}_{{\rm min}}$. Hence, parts (1.1.2) and (1.1.3) of Theorem 1.1 are sharp.
Example 4.12 ([BS19, Example 2.12]).
Following the notation in Example 4.11, we have $G_{\operatorname{ab}}=G=\mathbb{C}^{\times}$. Moreover, it is well-known that $\overline{\mathbf{O}}_{{\rm min}}$ is not $\mathbb{Q}$-factorial and admits a symplectic resolution by collapsing the zero section of $T^{*}\mathbb{P}^{n-1}$.
Now, let $n=2k$ and let $G^{\prime}\coloneqq\mathbb{C}^{\times}\rtimes\mathbb{Z}_{2}$, where $s\in\mathbb{Z}_{2}$ acts on $\mathbb{C}^{\times}$ by $s(\lambda)=\lambda^{-1}$. Then $G^{\prime},G^{\prime}]=\mathbb{C}^{\times}$ and $G_{\operatorname{ab}}\cong\mathbb{Z}_{2}$. Let $s$ act on $V_{k}=\mathbb{C}^{2k}$ as follows:
$s\cdot(x_{1},\dots,x_{k},x_{k+1},\dots,x_{2k})=(x_{k+1},\dots,x_{2k},x_{1},\dots,x_{k}).$
Then $V_{k}$ is a $2k$-large representation of $G^{\prime}$ and $T^{*}V_{k}/\!\!/\!\!/G^{\prime}$ is isomorphic to $\overline{\textbf{O}}_{\min}/\mathbb{Z}_{2}$, which is isomorphic to the nilpotent orbit closure $\overline{\textbf{O}}_{[2^{2},1^{2k-4}]}$ in $\mathfrak{sp}_{2k}$, or more precisely,
$\overline{\textbf{O}}_{[2^{2},1^{2k-4}]}=\left\{A\in\mathfrak{sp}_{2k}\mid A^{2}=0,\,\operatorname{rank}(A)\leq 2\right\}$
The nilpotent closure $\overline{\textbf{O}}_{[2^{2},1^{2k-4}]}$ is next to the minimal orbit closure in the sense that it is contained in all other nilpotent orbit closures, and $\operatorname{Cl}(\overline{\textbf{O}}_{[2^{2},1^{2k-4}]})\cong\mathbb{Z}_{2}$ by [Fu03, Corollary 2.7]. These examples show that we cannot remove the triviality and finiteness conditions on $G_{\operatorname{ab}}$ in Theorem 1.3. In this sense, our Theorem 1.3 is also optimal.
5. Differential operators on orbit spaces
Let $Z$ be an irreducible normal affine variety. Then the rings of differential operators on $Z$ and $Z_{\operatorname{reg}}$ coincide, i.e., $\mathcal{D}(Z)=\mathcal{D}(Z_{\operatorname{reg}})$. Let $S(Z)$ be the graded ring
$S(Z)\coloneqq\bigoplus_{p\geq 0}H^{0}(Z_{\operatorname{reg}},{\sf S}^{p}T_{Z_{\operatorname{reg}}})\cong\mathcal{O}(T^{*}Z_{\operatorname{reg}}).$
An element $\phi\in S^{k}(Z):=H^{0}(Z_{\operatorname{reg}},{\sf S}^{k}T_{Z_{\operatorname{reg}}})$ is called an order $k$ symbol. The ring $\mathcal{D}(Z)$ has a filtration $\{\mathcal{D}^{k}(Z)\}$ according to the order of differentiations, and the associated graded ring $\operatorname{gr}\mathcal{D}(Z)$ is commutative. Then the order $k$ symbol map gives a linear map $\sigma_{Z,k}\colon\mathcal{D}^{k}(Z)\rightarrow S^{k}(Z)$, which defines an injective homomorphism of graded algebras, called the graded symbol map:
$\bar{\sigma}_{Z}\colon\operatorname{gr}\mathcal{D}(Z)\longrightarrow S(Z).$
The symbol map is the map
$\sigma_{Z}\colon\mathcal{D}(Z)\longrightarrow S(Z),$
defined by $\sigma_{Z}(D)=\sigma_{Z,k}(D)$ if $D$ has order $k$. An order $k$ symbol $\phi$ quantizes into $D$ if $D\in\mathcal{D}^{k}(Z)$ has order $k$ and the principal symbol $\sigma_{Z,k}(D)$ of $D$ is $\phi$. If $Z$ is non-singular, then every symbol can be quantized, but for singular $Z$, the surjectivity of the symbol map is not well-understood.
We will consider the case where $Z$ is a quotient, i.e. Z=Y//GZ=Y/\!/G, where $Y$ is an irreducible non-singular affine $G$-variety. This problem was investigated in [Sch95]. Recall that there exists a canonical morphism $(\pi_{Y})_{*}\colon\mathcal{D}(Y)^{G}\rightarrow\mathcal{D}(Z)$ that preserves the filtrations by order, where $\pi_{Y}\colon Y\rightarrow Z$ is the canonical quotient morphism. Then $(\pi_{Y})_{*}$ is called graded surjective if the induced graded morphism $\operatorname{gr}(\pi_{Y})_{*}\colon\operatorname{gr}(\mathcal{D}(Y)^{G})\longrightarrow\operatorname{gr}(\mathcal{D}(Z))$ is surjective.
As an application of Theorem 1.1, we give a short proof of the graded surjectivity of $2$-large quotients established in [Sch95], and also show the surjectivity of the symbol map of such varieties.
Proof of Theorem 1.2.
The shell $N\subset T^{*}Y$ is an irreducible normal affine variety such that $\mathcal{O}(T^{*}Y)\rightarrow\mathcal{O}(N)$ is surjective. Since $G$ is reductive, taking invariant subrings yields a surjection
$S(Y)^{G}=\mathcal{O}(T^{*}Y)^{G}\longrightarrow\mathcal{O}(N)^{G}.$
and an isomorphism
$\operatorname{gr}(\mathcal{D}(Y)^{G})\longrightarrow(\operatorname{gr}\mathcal{D}(Y))^{G}.$
Since $Y$ is a non-singular affine variety, the graded symbol map $\bar{\sigma}_{Y}\colon\operatorname{gr}\mathcal{D}(Y)\rightarrow S(Y)$ is an isomorphism, so we get a surjection
[…]
Thanks to Theorem 1.1, there exists an canonical isomorphism
$\iota\colon\mathcal{O}(N)^{G}\cong\mathcal{O}(T^{*}Y/\!\!/\!\!/G)\longrightarrow S(Z_{\operatorname{reg}})=S(Z),$
and one easily checks that the following diagram is commutative:
[…]
Since $\alpha$ is surjective and $\iota$ is an isomorphism, it follows that $\bar{\sigma}_{Z}$ is surjective and hence an isomorphism, which implies that $\operatorname{gr}(\pi_{Y})_{*}$ is surjective. As $\mathcal{D}(Z)\to{\rm gr}\mathcal{D}(Z)$ is surjective, $\sigma_{Z}$ is also surjective. ∎
Appendix A Hamiltonian slice theorem
A.1. Set-up
Let $G$ be a complex reductive algebraic group with Lie algebra $\mathfrak{g}$ and let $X$ be an affine quasi-Hamiltonian $G$-variety. Let $x\in N\cap X_{\operatorname{reg}}$ be a given point such that the orbit $O_{x}=Gx$ is closed. Let $H$ be the isotropy subgroup of $G$ at $x$ with Lie algebra $\mathfrak{h}$. As the orbit $O_{x}=Gx$ is closed, $H$ is reductive by Matsushima’s theorem [Mat60], and it follows from Luna’s slice theorem (see [Lun73]) that there exists a locally closed non-singular affine $H$-stable subset $S\subset X$ containing $x$, called an étale slice, such that the morphism
$\varphi\colon Y\coloneqq G\times_{H}S\longrightarrow X,\quad[g,s]\longmapsto gs,$
is excellent and the image $GS\subset X$ is open affine and is contained in $X_{\rm reg}$ as $S$ is non-singular.
The symplectic form on $X_{\operatorname{reg}}$ can be pulled back to a symplectic form $\omega_{Y}$ on $Y$, making $(Y,\omega_{Y})$ a non-singular Hamiltonian $G$-variety, and the following diagram is commutative:
[…]
where $\mu_{X}=\mu_{G,X}$ and $\mu_{Y}=\mu_{G,Y}$. Let $y=[e,x]\in Y$. Then $\varphi(y)=x$.
Lemma A.1.
The orbit $O_{y}\coloneqq Gy$ is isotropic with respect to $\omega_{Y}$.
Proof.
By assumption, $O_{y}$ is contained in $\mu_{Y}^{-1}(0)$, so the restricted Hamiltonian function $H_{\xi}|_{O_{y}}$ is a constant and the restriction of the Hamiltonian vector field $V_{\xi}$ to $O_{y}$ is tangent to $O_{y}$ for any $\xi\in\mathfrak{g}$. In particular, we have
$\omega_{Y}(V_{\xi},V_{\xi^{\prime}})|_{O_{y}}=(dH_{\xi})|_{O_{y}}(V_{\xi^{\prime}}|_{O_{y}})=d(H_{\xi}|_{O_{y}})(V_{\xi^{\prime}}|_{O_{y}})=0.$
for any $\xi$, $\xi^{\prime}\in\mathfrak{g}$. Then the result follows since the $V_{\xi}$’s, $\xi\in\mathfrak{g}$, generate $T{O_{y}}$. ∎
A.2. Auxiliary constructions
Let us consider the following composition of linear maps:
[…]
The image of $\mathfrak{g}$ in $T_{y}Y$ is exactly $T_{y}O_{y}$. Let $\mathfrak{m}_{Y,y}\subset\mathcal{O}(Y)$ be the maximal ideal corresponding to $y$. Then $H_{\xi}\in\mathfrak{m}_{Y,y}$ for any $\xi\in\mathfrak{g}$ as $\mu_{Y}(y)=\mu_{X}(x)=0$. Moreover, under the canonical identification $T_{y}^{*}Y\cong\mathfrak{m}_{Y,y}/\mathfrak{m}_{Y,y}^{2}$, the linear map $\alpha$ is identified with the following map:
$\mathfrak{g}\longrightarrow\mathfrak{m}_{Y,y}/\mathfrak{m}_{Y,y}^{2},\quad\xi\longmapsto[H_{\xi}].$
Since $H$ is the isotropy subgroup and $\omega_{Y,y}$ is an isomorphism, the kernel of $\alpha$ is the subalgebra $\mathfrak{h}$. Since $H$ is reductive and $\mathfrak{h}$ is $H$-stable, we can choose an $H$-stable subspace $\mathfrak{l}$ of $\mathfrak{g}$ such that $\mathfrak{g}\cong\mathfrak{h}\oplus\mathfrak{l}$ and the induced linear map
$\alpha|_{\mathfrak{l}}\colon\mathfrak{l}\longrightarrow T_{y}Y\longrightarrow T_{y}^{*}Y$
is injective.
Let $S_{y}(\cong S)$ be the fiber of $Y=G\times_{H}S\rightarrow G/H$ over $[H]$. Then $y\in S_{y}$. Let $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ be the maximal ideal corresponding to $y$. Then we have a composition of linear maps
$\beta\colon\mathfrak{g}\longrightarrow T_{y}Y\longrightarrow T_{y}^{*}Y\cong\mathfrak{m}_{Y,y}/\mathfrak{m}_{Y,y}^{2}\longrightarrow T^{*}_{y}S_{y}\cong\mathfrak{m}_{S_{y},y}/\mathfrak{m}_{S_{y},y}^{2},$
which sends $\xi\in\mathfrak{g}$ to $H_{\xi}|_{S_{y}}]$.
Lemma A.2.
The linear map $\beta|_{\mathfrak{l}}\colon\mathfrak{l}\rightarrow T_{y}^{*}S_{y}$ is injective.
Proof.
Assume to the contrary that $\beta|_{\mathfrak{l}}$ is not injective. Then there exists a non-zero element $\xi\in\mathfrak{l}$ such that
$dH_{\xi}(y)(T_{y}S_{y})=\omega_{Y,y}(V_{\xi}(y),T_{y}S_{y})=0.$
On the other hand, since $0\not=V_{\xi}(y)\in T_{y}O_{y}$ and $O_{y}$ is isotropic with respect to $\omega_{Y,y}$ by Lemma A.1, one gets
$dH_{\xi}(y)(T_{y}O_{y})=\omega_{Y,y}(V_{\xi}(y),T_{y}O_{y})=0.$
In particular, as $T_{y}Y\cong T_{y}S_{y}\oplus T_{y}O_{y}$, one derives
$\omega_{Y,y}(V_{\xi}(y),T_{y}Y)=0,$
which is a contradiction because $\omega_{Y,y}$ is non-degenerate. ∎
Let $L$ be the subspace of $\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$ generated by $H_{\xi}$ with $\xi\in\mathfrak{l}$. By Lemma A.2, we have $L\cap\mathfrak{m}_{S_{y},y}^{2}=\{0\}$. In particular, since $L$ and $\mathfrak{m}^{2}_{S_{y},y}$ are $H$-stable subspaces of $\mathfrak{m}_{S_{y},y}$ and $H$ is reductive, we may choose a $H$-stable complement $F$ of $L\oplus\mathfrak{m}^{2}_{S_{y}}$ in $\mathfrak{m}_{S_{y},y}$. Then the natural inclusion
$F\oplus L\longrightarrow\mathfrak{m}_{S_{y},y}\subset\mathcal{O}(S_{y})$
induces a dominant morphism $\nu\colon S_{y}\rightarrow(F\oplus L)^{*}=F^{*}\oplus L^{*}$ and a linear isomorphism
$T_{y}^{*}S_{y}\cong\mathfrak{m}_{S_{y},y}/\mathfrak{m}_{S_{y},y}^{2}\cong F\oplus L.$
Clearly $\nu$ is étale at $y$ and maps the fixed point $y$ to $0\in F^{*}\oplus L^{*}$. Then Luna’s fundamental lemma [Lun75] implies that $\nu$ is an excellent morphism in an open affine $H$-saturated neighborhood of $y$. Thus, we may assume that $\nu$ itself is excellent after shrinking $S_{y}$. Now, let $U$ be the preimage $\nu^{-1}(F^{*})$. Then $U$ is non-singular since $\nu$ is étale.
Lemma A.3.
The variety $U$ is the fiber over $0\in\mathfrak{l}^{*}$ of the following natural composition:
[…]
Proof.
Note that $L$ is generated by the functions $H_{\xi}|_{S_{y}}$, $\xi\in\mathfrak{l}$. ∎
A.3. Proof of Theorem 3.2
Lemma A.4.
The restriction of $\omega_{Y}$ to $U$ is symplectic in an open neighborhood of $y$.
Proof.
By the openness of the non-degeneracy of $\omega_{Y}|_{U}$, it is enough to prove the non-degeneracy of $\omega_{Y}|_{U}$ at $y$.
Consider the linear map $\alpha|_{\mathfrak{l}}\colon\mathfrak{l}\rightarrow T_{y}Y\rightarrow T_{y}^{*}Y$. Then the image of $\mathfrak{l}$ in $T_{y}Y$ is exactly $T_{y}O_{y}$. Let us identify $\mathfrak{l}$ with its image in $T_{y}^{*}Y$; that is, the subspace generated by $dH_{\xi}(y)$ for $\xi\in\mathfrak{l}$. Let $\operatorname{Ann}(\mathfrak{l})\subset T_{y}Y$ be the annilhilator of $\mathfrak{l}$, which contains the tangent space $T_{y}U$ by the construction of $U$. In particular, since $T_{y}O_{y}$ is isotropic with respect to $\omega_{Y,y}$, a dimension count then yields
$\operatorname{Ann}(\mathfrak{l})=(T_{y}O_{y})^{\perp_{\omega_{Y,y}}}=T_{y}O_{y}\oplus T_{y}U.$
The result follows because $T_{y}O_{y}$ is isotropic with respect to $\omega_{Y,y}$ by Lemma A.1. ∎
Proof of Theorem 3.2.
According to Lemma A.4, after removing the degeneracy locus of $\omega_{Y}|_{U}$ in $U$ if necessary, we may assume that $U\subset Y$ is a non-singular affine Hamiltonian $H$-variety such that the following diagram commutes:
[…]
The closed subscheme $S_{y}\cap N_{Y}$ of $S_{y}$ is defined by the functions $H_{\xi}|_{S_{y}}$, $\xi\in\mathfrak{g}$. On the other hand, $U$ is defined by the functions $H_{\xi}|_{S_{y}}$, $\xi\in\mathfrak{l}$ and the commutative diagram above says that $N_{U}$ is defined by the functions $H_{\xi}|_{U}$, $\xi\in\mathfrak{h}$. As $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{l}$, the natural inclusion
$N_{U}^{\operatorname{red}}\longrightarrow(S_{y}\cap N_{Y})^{\operatorname{red}}$
is actually an open embedding. Finally the statement follows from the fact that the restriction $\varphi|_{\varphi^{-1}(Z)}\colon\varphi^{-1}(Z)\rightarrow Z$ to any $G$-invariant closed subset $Z\subset X$ is again excellent. ∎
The following result is a by-product of the proof of Theorem 3.2.
Corollary A.5.
In Theorem 3.2, there exists an étale slice $S$ containing $U$, a $G_{x}$-stable decomposition
$T_{x}X=\mathfrak{g}/\mathfrak{g}_{x}\oplus T_{x}U\oplus(\mathfrak{g}/\mathfrak{g}_{x})^{*}$
and a commutative diagram with excellent $G_{x}$-equivariant horizontal maps:
[…]
Proof.
Since $\varphi$ is a $G$-equivariant étale morphism, it follows from the paragraph before Lemma A.3 that there exists a $G_{x}$-stable decomposition
$T_{x}X\cong T_{x}O_{x}\oplus F^{*}\oplus L^{*},$
such that $L^{*}\cong\mathfrak{l}^{*}$ and $F^{*}\cong T_{x}U$ as $G_{x}$-modules. So the statement follows from the fact that $T_{x}O_{x}$ and $\mathfrak{l}$ are isomorphic to $\mathfrak{g}/\mathfrak{g}_{x}$ as $G_{x}$-modules. ∎